OpenStudy (anonymous):

Find the nth partial sum for a_n = 1/(n+1) - 1/(n+2)

OpenStudy (anonymous):

put n=0, 1/1-1/2 put n=1, 1/2 -1/3 put n=2, 1/3-1/4 : : : nth term = 1/(n+1) - 1/(n+2) sum of n terms = (1/1-1/2)+(1/2 -1/3)+(1/3-1/4)+...+1/(n+1) - 1/(n+2) ^ ^ ^ ^ |______| |______| cancelled cancelled =1-1/(n+2)